OpenStudy (anonymous):

Solve for the variable- x+6-x+4=10-x What is the answer and HOW DO YOU do problems like this???? Steps please

OpenStudy (bahrom7893):

x=0

OpenStudy (anonymous):

x=0

OpenStudy (bahrom7893):

x+6-x+4=10-x x-x+6+4=10-x 0+10=10-x 10=10-x 10-10=10-x-10 0=-x x=0